The value of Spearman's rank correlation coefficient of a set of non-repeating values was found to be 2/3. The sum of the squares of difference between the corresponding ranks was 55. Find the number of pairs.


1
Expert's answer
2022-06-13T16:24:04-0400

"\rho=1-\frac{6\Sigma d_i^2}{n(n^2-1)}"

"\frac{2}{3}=1-\frac{6*55}{n(n^2-1)}"

"n(n^2-1)=990"

n=10